40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 3 19:09:56 2008, Yoichi, DAQ, Computers, c1iovme power cycled 
    Reply  Thu Sep 4 00:38:14 2008, rana, Update, PSL, c1iovme power cycled 
       Reply  Thu Sep 4 07:29:52 2008, Yoichi, Update, PSL, c1iovme power cycled 
Message ID: 917     Entry time: Wed Sep 3 19:09:56 2008     Reply to this: 918
Author: Yoichi 
Type: DAQ 
Category: Computers 
Subject: c1iovme power cycled 
When I tried to measure the sideband power of the FSS using the scan of the reference cavity, I noticed that the RC trans. PD signal was not
properly recorded by the frame builder.
Joe restarted c1iovme software wise. The medm screen said c1iovme is running fine, and actually some values were recorded by the FB.
Nonetheless, I couldn't see flashes of the RC when I scanned the laser frequency.
I ended up power cycling the c1iovme and run the restart script again. Now the signals recorded by c1iovme look fine.
Probably, the DAQ boards were not properly initialized only by the software reset.
I will re-try the sideband measurement tomorrow morning.
ELOG V3.1.3-